# Prediction of Cellular Malignancy Using Electrical Impedance Signatures and Supervised Machine Learning

### v2

@@ Abstract
Bioelectrical properties of cells such as relative permittivity, conductivity, and characteristic time constants vary significantly between healthy and malignant cells across different frequencies. These distinctions provide a promising foundation for diagnostic and classification applications. This study systematically reviewed 33 scholarly articles to compile datasets of quantitative bioelectric parameters and evaluated their utility in predictive modeling. Three supervised machine learning algorithms- Random Forest (RF), Support Vector Machine (SVM), and K-Nearest Neighbor (KNN) were implemented and tuned using key hyperparameters to assess classification performance. Model effectiveness was evaluated using accuracy and F1 score as performance metrics. Results demonstrate that Random Forest achieved the highest predictive accuracy of   90% when configured with a maximum depth of 4 and 100 estimators. These findings highlight the potential of integrating bioelectrical property analysis with machine learning for improved diagnostic decision-making. Similarly, for KNN and SVM, the F1 score peaked at approximately 78% and 76.5%, respectively. Future work will explore incorporating additional discriminative features, leveraging stimulated datasets, and optimizing hyperparameter through advanced search strategies. Ultimately, hardware prototype with embedded micro-electrodes and real-time control systems could pave the path for practical diagnostic tools capable of in-situ cell classification.
@@ 1 Introduction
According to the Centers for Disease Control and Prevention (CDC) and the National Cancer Institute, approximately 614,000 cancer-related deaths were reported in the United States in 2023. Additionally, an estimated 1.9 million and 2.0 million new cancer cases were reported in 2022 and 2025, respectively. Several factors influence cancer prognosis including cancer type, (ii) stage at diagnosis, (iii) cellular characteristics, and (iv) histological grade, among others [1]. Notably, changes in the electrical properties and biochemical composition of cells have emerged as significant indicators for disease diagnosis and progression [2]. Early detection is strongly correlated with improved survival outcomes, underscoring the importance of timely diagnosis and monitoring to reduce mortality and expand treatment options [3],[4],[5].
@@
One of the key advantages of electrical characterization of cells is its non-invasive and label-free nature [6]. Unlike fluorescence -based techniques, dielectric measurement does not require staining or labeling, which can potentially alter cellular behavior. Moreover, cells remain viable after measurement, enabling further downstream analysis or long-term studies if needed. Another notable advantage, driven by advancements in medical technology, is that real-time dynamic electrical measurements enable continuous cell monitoring and facilitate high-throughput screening.
@@
Cheung et al. (2005) developed a microfabricated impedance spectroscopy flow cytometer for the rapid dielectric characterization of cells across a range of frequencies [7]. They utilized electrical features such as amplitude, phase, and opacity to actively discriminate between cell types without the use of molecular markers. Although the study was limited to red blood cells (RBC), the approach holds potential for broader applications, including the measurements of cell conductivity and capacitance for early-stage apoptosis or cancer diagnostics.
@@
Similarly, Gawad et al. (2001) introduced a cytological tool capable of high-speed cell counting and separation, operating at a sampling rate of 100 samples per second [8]. Their chip-based flow cytometer employed a differential pair of microelectrodes to measure impedance over a frequency range of 100 kHz to 15 MHz. This system can be extended to distinguish between normal and malignant cells based on differences in bioimpedance or conductivity.
@@
Our previous works focused on electrical impedance and transmembrane potential as biophysical indicators for cancer cell diagnosis [9],[2]. The study revealed that cancer cells exhibit higher electrical conductivity and greater dielectric loss compared to their healthier counterparts. Additionally, cancer cells tend to be depolarized, often displaying elevated intracellular sodium ion concentrations despite similar external ionic conditions [10],[11],[12]. For instance, ovarian tumor cells have reported to possess a transmembrane potential of approximately -5 mV, while MCF -7 breast cancer cells range between -37 to -38 mV , and MDA- MB-468 cells exhibit a potential around -30 mV [13]. In our previous works, we also investigated that both the bioimpedance and optical properties of breast cancer cells can be used to enhance the precision of malignant cell identification [2],[9],[14]. Electrical measurements were conducted in the microwave frequency range, while the optical characteristics of normal and cancer cells were evaluated in the visible and near- infrared spectral (NIR) regions. Similarly, Alfano et al. conducted several pioneering studies utilizing optical techniques to distinguish precancerous and cancerous tissues based on their characteristic absorption wavelength [15],[16],[17]. However, despite their potential, optical diagnostic techniques face significant challenges, including :(i) limited spatial resolution, (ii) prolonged acquisition time, and (iii) reduced sensitivity in complex biological environments [18].
@@
Measurable bioelectrical differences between cancerous and non-cancerous cells provide a valuable feature space for computational classification. By leveraging machine learning algorithms, such as Support Vector Machine (SVM), Random Forest (RF), Deep Neural Network (DNN) or K-Nearest Neighbor (KNN), it is possible to learn patterns in electrical impedance, dielectric loss and transmembrane potential that distinguish malignant cells from healthy ones [19],[20],[21],[22],[23]. These models can be trained on labeled datasets derived from impedance cytometry or microelectrode array recordings, enabling real time non-invasive cancer diagnostics with high accuracy. Moreover, the ability of machine learning to capture complex, non-linear relationships enhance the diagnostic potential beyond traditional threshold-based methods. Fig.1 shows the schematic of the sequence of steps for the proposed model discussed earlier. Bioelectrical measurements -such as impedance, conductivity, and dielectric loss -are acquired from test samples using microelectrodes. These features are then input into a pretrained machine learning classifier for binary classification, distinguishing between malignant and healthy cells based on their electrical signatures.
@@
This study investigated and compared the performance of three supervised machine learning models- Random Forest (RF), Support Vector Machine (SVM) and K-Nearest Neighbor (KNN)- for classifying cell samples based on their bioimpedance properties. Labelled datasets were compiled from 33 scholarly journals, and model hyperparameters were systematically tuned to enhance predictive accuracy. Performance was evaluated using accuracy and F1 score, with the model achieving the highest metric proposed as a potential reference standard for malignancy identification.
@@ 2.1 Bioelectrical Properties and Cole-Cole plot
Cancer progression is accompanied by distinct biophysical and biochemical changes at the cellular level, many of which affect the structural and electrical properties of the cell membrane and cytoplasm [2],[9]. As a result electrical characterization techniques have emerged as a valuable tool in cancer research, offering non-invasive label-free methods to distinguish between normal, premalignant and malignant cells. Among these techniques, dielectrophoresis (DEP) has shown promise in detecting subtle differences in detecting dielectric properties such as permittivity, conductivity and membrane capacitance [24],[25],[26],[27],[2],[28]. These properties which are frequency dependent can be modeled using frameworks like the Cole-Cole equation to reveal patterns that correlate with malignancy. The ability to extract quantitative dielectric signatures makes the approaches particularly suitable for early detection, classification, and monitoring of cancer progression in vitro and in vivo settings.
@@
Mathematically, the internal electric field ($E_{in}$) and surface potential ($\Phi$) of a homogeneous dielectric sphere subjected to a uniform external electric field are given by equations (1) and (2): [29],[30]:
@@
A biological cell can be electrically modeled as a combination of resistive and capacitive elements [31],[32],[33]. The cytoplasm, rich in ions, is represented by an internal resistance, Ri reflecting its conductive properties. In contrast, the cell membrane, primarily composed of a lipid bilayer, acts as a dielectric barrier and is modeled as a parallel combination of resistance and capacitance. This RC configuration accounts for the membrane’s ability to store and partially block charge flow. Importantly, the overall importance of the cell is frequency-dependent: at low frequencies, the capacitive nature of the membrane dominates, impending current flow; while at higher frequencies, the capacitive reactance decreases, allowing current to pass through resistive pathways more easily [34],[35],[36].
Iqbal et al. (2019) presented an equivalent electrical model of an isolated single biological cell composed of resistive and capacitive elements [37]. In this model, the cell is represented by distinct compartments corresponding to the cytoplasm, nucleus and cell membrane. The capacitive components arise primarily from the charge-separation layers such as cell membrane and nuclear envelope, while the cytoplasm and nucleus are modeled as lossy conductive media and therefore contribute predominantly resistive components.
@@
Giannoukis and Min (2014) explored the mathematical and physical modeling of the dynamic electrical bioimpedance of cells [38]. Fig. 3(a) illustrates the Frickle-Morse equivalent circuit model representing a biological cell. Fig.3 (b) and (c) demonstrate the variation in electric field distribution within normal and cancerous cells, representing how differences in dielectric properties influence the field patterns. The increased permittivity of cancerous cells alters the dielectric response, resulting in more concentrated electric field distribution compared to non-tumorous.
@@
Fig. 3(d) illustrates the schematic of an impedance-based cell analysis technique, in which a small AC excitation current is applied across microelectrodes immersed in a conductive saline medium containing suspended cells. The presence of suspended cells perturbs the resulting current flow due to contrasts in their frequency-dependent dielectric properties, enabling the extraction of electrical impedance in a label free and real-time manner. This approach has been widely employed for dynamic cell analysis and characterization [37],[39],[40],[41]. Under the assumption of a dilute cell suspension and a simplified single-shell electrical representation, the effective impedance of the cell-electrolyte system can be approximated using equation (4)[39].
@@
The frequency-depending dielectric contrast between the cell membrane and intracellular component leads to a measurable variation in current flow, from which the impedance spectrum can be extracted [41]. In Electric cell-substrate impedance sensing (ECIS), impedance is defined as the ratio of the measured voltage to the applied current across the electrodes over a range of AC frequencies[42]. Fig. 3 (e) presents a complex equivalent circuit model of a biological cell inspired by the Fricke-Morse single-shell model. The cytoplasm, owing to its intracellular conductivity is represented by an intracellular resistance $R_{i}$ while both the cell membrane and nucleus are modeled using parallel resistive and capacitive elements, characterized by membrane capacitance $C_{m}$ and nuclear capacitance $C_{n}$ , respectively. These capacitive components govern the frequency-dependent electrical behavior of the cell through capacitive coupling, as illustrated in Fig. 3(f). The capacitive reactance of the cell membrane is given by equation (5):
@@
Several impedance-based techniques have been developed to protect the electrical properties of biological cells, each differing in electrode configuration, frequency range, sensitivity, and application domain. These methods exploit the frequency-dependent dielectric response of cell membranes and intracellular components to extract distinctive electrical impedance signatures associated with cellular morphology and physiological state. Table LABEL:tab:model_comparison1summarizes representative impedance measurements techniques reported in the literature, highlighting the cell types studied, key electrical parameters extracted, and their reported applications in tumor cell characterization.
@@
Electrical Impedance Spectroscopy (EIS) Analyzes complex impedance spectrum using equivalent circuit models over a broad range. Label-free characterization of dielectric properties and cell discrimination. SW403, Jurkat, and THP-1 10 Hz–1 MHz $R_{ct}$ (SW403/Jurkat): 2250 $\Omega$; THP-1: 2000 $\Omega$ [46]
@@
Microfluidic cell impedance analysis Measures voltage/impedance changes as single cells flow through microchannels under AC excitation. Label-free discrimination based on size, membrane, and conductivity. A549 tumor cells, RBC, and WBC 0.1–10 MHz A549: 4.2–79.9 mV; RBC: 2.4–19 mV; WBC: 1.7–10.5 mV [47]
The Cole- Cole equation is the relaxation model that is used to describe the dielectric relaxation, and the model can also be adapted to compare the separation in dielectric properties of cancerous and non-cancerous breast cancer cells [48]. Fig. 3 shows the Cole-Cole plot illustrating the separation in dielectric properties between cancerous and non-tumorous breast cell lines. The plot generated from dielectric spectroscopy data of various breast cell lines (including both cancerous and non-tumorigenic types), demonstrates a clear clustering pattern. The spatial distribution of the complex permittivity values suggests the existence of a separating hyperplane, enabling potential binary classification based on the dielectric properties. Further quantitative analysis (e.g. linear discrimination analysis or SVM classification) is warranted to validate the robustness and generalizability of this observed separation and is the focus of the following section.
Figure 4: Cole-Cole plot illustrating the separation in dielectric properties between cancerous and non-tumorous breast cell lines.
@@
Similar studies by Roberts et al. (2005) investigated the progression of mouse ovarian surface epithelial (MOSE) cells [49]. This work analyzed cellular and molecular changes across the different stages, from pre-malignant, non-tumorigenic states to highly aggressive malignant phenotypes. The model includes four transitional states including MOSE-E (early preneoplastic stage) and MOSE-L (the most aggressive stage) with high tumorigenic potential.
@@
In addition, to dielectric spectroscopy, studies on dielectrophoresis (DEP) – a phenomenon in which a force is exerted on dielectric particle in a non-uniform electric field has also been employed to investigate the different stages of MOSE cells [50]. The results indicate that as malignancy progresses the membrane capacitance decreases from approximately 26 mF/m2 to 15 mF/m2. This decline highlights the difference in dielectric properties and membrane composition between malignant and non-tumorous cells, reinforcing the potential of electrical characterization methods in distinguishing cancer progression.
@@
While dielectric properties provide a promising biophysical basis for distinguishing between malignant and non-malignant cells, the complexity and variability of the measured parameters necessitate robust computational tools for accurate classification. Artificial Intelligence (AI) and Machine Learning (ML) algorithms have been increasingly adopted to analyze bioelectrical data. These models can learn patterns from features such as membrane capacitance, conductivity, and impedance spectra to reliably differentiate between healthy and cancerous cells. Among various ML techniques, models such as Random Forest (RF), Support Vector Machine (SVM) and K-Nearest Neighbor (KNN), and Neural Network have shown promising results in biomedical classification tasks due to their ability to handle non-linear data and high dimension feature space [51],[52],[53],[54],[55],[56],[9].

### v3

@@ Abstract
Bioelectrical properties of cells such as relative permittivity, conductivity, and characteristic time constants vary significantly between healthy and malignant cells across different frequencies. These distinctions provide a promising foundation for diagnostic and classification applications. This study systematically reviewed 20 scholarly articles to compile 535 datasets of quantitative bioelectric parameters and evaluated their utility in predictive modeling. Three supervised machine learning algorithms- Random Forest (RF), Support Vector Machine (SVM), and K-Nearest Neighbor (KNN) were implemented and tuned using key hyperparameters to assess classification performance. Model effectiveness was evaluated using accuracy and F1 score as performance metrics. Results demonstrate that Random Forest achieved the highest predictive accuracy of   90% when configured with a maximum depth of 4 and 100 estimators. These findings highlight the potential of integrating bioelectrical property analysis with machine learning for improved diagnostic decision-making. Similarly, for KNN and SVM, the F1 score peaked at approximately 78% and 76.5%, respectively. Future work will explore incorporating additional discriminative features, leveraging stimulated datasets, and optimizing hyperparameter through advanced search strategies. Ultimately, hardware prototype with embedded micro-electrodes and real-time control systems could pave the path for practical diagnostic tools capable of in-situ cell classification.
@@ 1 Introduction
According to the Centers for Disease Control and Prevention (CDC) and the National Cancer Institute, approximately 614,000 cancer-related deaths were reported in the United States in 2023. Additionally, an estimated 1.9 million and 2.0 million new cancer cases were reported in 2022 and 2025, respectively. Several factors influence cancer prognosis including cancer type, (ii) stage at diagnosis, (iii) cellular characteristics, and (iv) histological grade, among others [Rakha2010]. Notably, changes in the electrical properties and biochemical composition of cells have emerged as significant indicators for disease diagnosis and progression [Hossain2020]. Early detection is strongly correlated with improved survival outcomes, underscoring the importance of timely diagnosis and monitoring to reduce mortality and expand treatment options [McPhail2015],[Muller2018],[vandenBergh2015].
@@
One of the key advantages of electrical characterization of cells is its non-invasive and label-free nature [NgocLe2019]. Unlike fluorescence -based techniques, dielectric measurement does not require staining or labeling, which can potentially alter cellular behavior. Moreover, cells remain viable after measurement, enabling further downstream analysis or long-term studies if needed. Another notable advantage, driven by advancements in medical technology, is that real-time dynamic electrical measurements enable continuous cell monitoring and facilitate high-throughput screening.
@@
Cheung et al. (2005) developed a microfabricated impedance spectroscopy flow cytometer for the rapid dielectric characterization of cells across a range of frequencies [Cheung2005]. They utilized electrical features such as amplitude, phase, and opacity to actively discriminate between cell types without the use of molecular markers. Although the study was limited to red blood cells (RBC), the approach holds potential for broader applications, including the measurements of cell conductivity and capacitance for early-stage apoptosis or cancer diagnostics.
@@
Similarly, Gawad et al. (2001) introduced a cytological tool capable of high-speed cell counting and separation, operating at a sampling rate of 100 samples per second [Gawad2001]. Their chip-based flow cytometer employed a differential pair of microelectrodes to measure impedance over a frequency range of 100 kHz to 15 MHz. This system can be extended to distinguish between normal and malignant cells based on differences in bioimpedance or conductivity.
@@
Our previous works focused on electrical impedance and transmembrane potential as biophysical indicators for cancer cell diagnosis [Hossain2021],[Hossain2020]. The study revealed that cancer cells exhibit higher electrical conductivity and greater dielectric loss compared to their healthier counterparts. Additionally, cancer cells tend to be depolarized, often displaying elevated intracellular sodium ion concentrations despite similar external ionic conditions [Kunzelmann2005],[Bortner2014],[Binggeli1986]. For instance, ovarian tumor cells have reported to possess a transmembrane potential of approximately -5 mV, while MCF -7 breast cancer cells range between -37 to -38 mV , and MDA- MB-468 cells exhibit a potential around -30 mV [Mahmoud2023]. In our previous works, we also investigated that both the bioimpedance and optical properties of breast cancer cells can be used to enhance the precision of malignant cell identification [Hossain2020],[Hossain2021],[HossainHossain2021]. Electrical measurements were conducted in the microwave frequency range, while the optical characteristics of normal and cancer cells were evaluated in the visible and near- infrared spectral (NIR) regions. Similarly, Alfano et al. conducted several pioneering studies utilizing optical techniques to distinguish precancerous and cancerous tissues based on their characteristic absorption wavelength [Alfano2016],[Alfano2013],[Alfano2005]. However, despite their potential, optical diagnostic techniques face significant challenges, including :(i) limited spatial resolution, (ii) prolonged acquisition time, and (iii) reduced sensitivity in complex biological environments [Licha2005]. While numerous studies have demonstrated the feasibility of utilizing raw impedance spectra or optical signatures for cancer detection, most existing approaches rely primarily on directly extracted measurement features, which often lack clear physical interpretability and generalizability across experimental platforms. To address this limitation, there is a growing need for frameworks that integrate electrical measurements with biophysical models to derive intrinsic cellular properties-such as membrane permittivity, cytoplasmic conductivity and transmembrane potential that more directly reflect underlying cellular physiology. Such physics guided representation can provide robust, interpretable biomarkers and form a stronger foundation for data driven learning algorithm.
@@
Measurable bioelectrical differences between cancerous and non-cancerous cells provide a valuable feature space for computational classification. When combined with equivalent circuit model and electromagnetic analysis, these measurements can be transformed into physically meaningful dielectric and conductive parameter.By leveraging machine learning algorithms, such as Support Vector Machine (SVM), Random Forest (RF), Deep Neural Network (DNN) or K-Nearest Neighbor (KNN), it is possible to learn patterns in electrical impedance, dielectric loss and transmembrane potential that distinguish malignant cells from healthy ones [Biau2016],[Rigatti2017],[Guo2003],[Hossain2024Authorea],[Suthaharan2016]. These models can be trained on labeled datasets derived from impedance cytometry or microelectrode array recordings, enabling real time non-invasive cancer diagnostics with high accuracy. Moreover, the ability of machine learning to capture complex, non-linear relationships enhance the diagnostic potential beyond traditional threshold-based methods. Fig.1 shows the schematic of the sequence of steps for the proposed model discussed earlier. Bioelectrical measurements -such as impedance, conductivity, and dielectric loss -are acquired from test samples using microelectrodes. These features are then input into a pretrained machine learning classifier for binary classification, distinguishing between malignant and healthy cells based on their electrical signatures.
@@
This study investigated and compared the performance of three supervised machine learning models- Random Forest (RF), Support Vector Machine (SVM) and K-Nearest Neighbor (KNN)- for classifying cell samples based on their bioimpedance properties. Labelled datasets were compiled from 20 scholarly journals with 535 datasets, and model hyperparameters were systematically tuned to enhance predictive accuracy. Performance was evaluated using accuracy and F1 score, with the model achieving the highest metric proposed as a potential reference standard for malignancy identification.
@@ 2.1 Bioelectrical Properties and Cole-Cole plot
Cancer progression is accompanied by distinct biophysical and biochemical changes at the cellular level, many of which affect the structural and electrical properties of the cell membrane and cytoplasm [Hossain2020],[Hossain2021]. As a result electrical characterization techniques have emerged as a valuable tool in cancer research, offering non-invasive label-free methods to distinguish between normal, premalignant and malignant cells. Among these techniques, dielectrophoresis (DEP) has shown promise in detecting subtle differences in detecting dielectric properties such as permittivity, conductivity and membrane capacitance [Pethig2010],[Pohl1971],[Nakano2013],[Hossain2023],[Hossain2020],[Lapizco2007]. These properties which are frequency dependent can be modeled using frameworks like the Cole-Cole equation to reveal patterns that correlate with malignancy. The ability to extract quantitative dielectric signatures makes the approaches particularly suitable for early detection, classification, and monitoring of cancer progression in vitro and in vivo settings.
@@
Mathematically, the internal electric field ($E_{in}$) and surface potential ($\Phi$) of a homogeneous dielectric sphere subjected to a uniform external electric field are given by equations (1) and (2): [Ref65],[Ref66]:
@@
A biological cell can be electrically modeled as a combination of resistive and capacitive elements [Brosseau2021],[Gowrishankar2003],[Ellappan2005]. The cytoplasm, rich in ions, is represented by an internal resistance, Ri reflecting its conductive properties. In contrast, the cell membrane, primarily composed of a lipid bilayer, acts as a dielectric barrier and is modeled as a parallel combination of resistance and capacitance. This RC configuration accounts for the membrane’s ability to store and partially block charge flow. Importantly, the overall importance of the cell is frequency-dependent: at low frequencies, the capacitive nature of the membrane dominates, impending current flow; while at higher frequencies, the capacitive reactance decreases, allowing current to pass through resistive pathways more easily [Trainito2015],[Mahesh2020],[Karmakar2024].
Iqbal et al. (2019) presented an equivalent electrical model of an isolated single biological cell composed of resistive and capacitive elements [Iqbal2019]. In this model, the cell is represented by distinct compartments corresponding to the cytoplasm, nucleus and cell membrane. The capacitive components arise primarily from the charge-separation layers such as cell membrane and nuclear envelope, while the cytoplasm and nucleus are modeled as lossy conductive media and therefore contribute predominantly resistive components.
@@
Giannoukis and Min (2014) explored the mathematical and physical modeling of the dynamic electrical bioimpedance of cells [Giannoukos2014]. Fig. 3(a) illustrates the Frickle-Morse equivalent circuit model representing a biological cell. Fig.3 (b) and (c) demonstrate the variation in electric field distribution within normal and cancerous cells, representing how differences in dielectric properties influence the field patterns. The increased permittivity of cancerous cells alters the dielectric response, resulting in more concentrated electric field distribution compared to non-tumorous.
@@
Fig. 3(d) illustrates the schematic of an impedance-based cell analysis technique, in which a small AC excitation current is applied across microelectrodes immersed in a conductive saline medium containing suspended cells. The presence of suspended cells perturbs the resulting current flow due to contrasts in their frequency-dependent dielectric properties, enabling the extraction of electrical impedance in a label free and real-time manner. This approach has been widely employed for dynamic cell analysis and characterization [Iqbal2019],[Ref68],[Ref69],[Ref70]. Under the assumption of a dilute cell suspension and a simplified single-shell electrical representation, the effective impedance of the cell-electrolyte system can be approximated using equation (4)[Ref68].
@@
The frequency-depending dielectric contrast between the cell membrane and intracellular component leads to a measurable variation in current flow, from which the impedance spectrum can be extracted [Ref70]. In Electric cell-substrate impedance sensing (ECIS), impedance is defined as the ratio of the measured voltage to the applied current across the electrodes over a range of AC frequencies[Ref71]. Fig. 3 (e) presents a complex equivalent circuit model of a biological cell inspired by the Fricke-Morse single-shell model. The cytoplasm, owing to its intracellular conductivity is represented by an intracellular resistance $R_{i}$ while both the cell membrane and nucleus are modeled using parallel resistive and capacitive elements, characterized by membrane capacitance $C_{m}$ and nuclear capacitance $C_{n}$ , respectively. These capacitive components govern the frequency-dependent electrical behavior of the cell through capacitive coupling, as illustrated in Fig. 3(f). The capacitive reactance of the cell membrane is given by equation (5):
@@
Several impedance-based techniques have been developed to protect the electrical properties of biological cells, each differing in electrode configuration, frequency range, sensitivity, and application domain. These methods exploit the frequency-dependent dielectric response of cell membranes and intracellular components to extract distinctive electrical impedance signatures associated with cellular morphology and physiological state. Table 1summarizes representative impedance measurements techniques reported in the literature, highlighting the cell types studied, key electrical parameters extracted, and their reported applications in tumor cell characterization.
@@
Electrical Impedance Spectroscopy (EIS) Analyzes complex impedance spectrum using equivalent circuit models over a broad range. Label-free characterization of dielectric properties and cell discrimination. SW403, Jurkat, and THP-1 10 Hz–1 MHz $R_{ct}$ (SW403/Jurkat): 2250 $\Omega$; THP-1: 2000 $\Omega$ [Ref75]
@@
Microfluidic cell impedance analysis Measures voltage/impedance changes as single cells flow through microchannels under AC excitation. Label-free discrimination based on size, membrane, and conductivity. A549 tumor cells, RBC, and WBC 0.1–10 MHz A549: 4.2–79.9 mV; RBC: 2.4–19 mV; WBC: 1.7–10.5 mV [Ref76]
Table2 compares electrical impedance parameters, including resistive and capacitive elements, reported for healthy and cancerous cells across different impedance-based measurement techniques and frequency ranges. These parameters play a critical role in characterizing cellular electrical behavior for diagnostic applications. A consistent trend observed across multiple studies is that healthy cells generally exhibit higher membrane capacitance compared to cancerous cells, with capacitance decreasing progressively as cellular invasiveness increases. This reduction is commonly attributed to alterations in membrane morphology, composition, and surface roughness associated with malignant transformation. Similarly, the magnitude of cellular resistance decreases in cancerous cells, reflecting increased ionic conductivity and membrane permeability resulting from disputed cellular architecture. Together, these impedance variations provide discriminative electrical signatures that can be exploited for cancer detection and phenotypic classification.
The frequency-dependent dielectric behavior of biological tissues and cells is commonly described using the complex permittivity formalism:
$\epsilon^{*}(\omega)=\epsilon^{\prime}(\omega)-j\epsilon^{\prime\prime}(\omega)$ (6)
where $\epsilon^{\prime}$ represents the real (energy storage) component and $\epsilon^{\prime\prime}$ denotes the imaginary (loss) component of permittivity.
To model dielectric relaxation in heterogeneous biological media, the Cole-Cole dispersion model is employed:
$\epsilon^{*}(\omega)=\epsilon_{\infty}+\frac{\epsilon_{s}-\epsilon_{\infty}}{1+(j\omega\tau)^{1-\alpha}}+\frac{\sigma_{i}}{j\omega\epsilon_{0}}$ (7)
where $\epsilon_{s}$ and $\epsilon_{\infty}$ denote the static and high-frequency permittivity limits, respectively, $\tau$ is the characteristic relaxation time, $\alpha$ ($0\leq\alpha\leq 1$) describes the distribution of relaxation times, $\sigma_{i}$ represents the ionic conductivity, and $\epsilon_{0}$ is the permittivity of free space.
For $\alpha=0$, the model reduces to classical Debye relaxation. Nonzero $\alpha$ values account for dispersion broadening associated with structural heterogeneity, membrane polarization, and intracellular complexity. In biological tissues, variations in water content, membrane integrity, and ionic composition significantly influence these parameters, leading to distinguishable trajectories in the complex permittivity plane [Ref85, Ref86].
The Cole-Cole representation therefore provides a physically grounded framework for comparing dielectric signatures for malignant and non-tumorigenic breast cell lines, enabling visualization of dispersion characteristics and subsequent quantitative classification analysis. The Cole-Cole relaxation model provides a robust framework for characterizing dielectric dispersion in biological tissues [Cole1941]. Fig. 4 illustrates the complex permittivity distribution of malignant (T-47D, MCF-7, MDA-MB-231) and non-tumorigenic MCF-10A breast epithelial cell lines, alongside reported tissue baselines [Ref82, Ref83, Ref84]. The theoretical Cole-Cole relaxation loci (gray arcs) are superimposed to conceptualize the dispersion behavior.
A distinct clustering pattern is observed in the complex permittivity plane in Fig. 4: low-water content fatty tissues occupy the lower dielectric regimes, whereas malignant cell lines consistently exhibit elevated real permittivity values ($\epsilon^{\prime}$) across the investigated frequency range. Importantly, the tumor-fat dielectric contrast persists into the millimeter-wave regime, supporting the broadband viability of electromagnetic diagnostic techniques.
The spatial distribution of the dielectric signatures visually suggests potential separability in the complex permittivity domain. Quantitative classifier analysis (e.g., linear discrimination analysis and Support Vector Machines) is performed in the following section to rigorously evaluate discriminative performance.
Table 2: Summary of reported resistive and capacitive electrical parameters of healthy and cancerous cell types from the literature.
Cell Type Nature Frequency Electrical Parameter Value/Units Ref.
HT29 + 10% SW48 Cancerous 20–200 kHz Impedance 1100–200 $\Omega$ [Ref77]
HT29 + 10% SW48 Cancerous 20–200 kHz Impedance 1460–300 $\Omega$ [Ref77]
[…]
MDCK II cells Healthy 500 Hz–64 kHz Capacitance 3.9–0.66 nF [Ref78]
[…]
AML 2 cells Cancerous 20–400 kHz Specific membrane capacitance 12 $\pm$ 1.44 mF/m2 [Ref81]
HL-60 cells Cancerous 20–400 kHz Specific membrane capacitance 14.5 $\pm$ 1.75 mF/m2 [Ref81]
Figure 4: Cole-Cole complex permittivity mapping of malignant and non-tumorigenic breast cell lines.
@@
Similar studies by Roberts et al. (2005) investigated the progression of mouse ovarian surface epithelial (MOSE) cells [Roberts2005]. This work analyzed cellular and molecular changes across the different stages, from pre-malignant, non-tumorigenic states to highly aggressive malignant phenotypes. The model includes four transitional states including MOSE-E (early preneoplastic stage) and MOSE-L (the most aggressive stage) with high tumorigenic potential.
@@
In addition, to dielectric spectroscopy, studies on dielectrophoresis (DEP) – a phenomenon in which a force is exerted on dielectric particle in a non-uniform electric field has also been employed to investigate the different stages of MOSE cells [Trainito2019]. The results indicate that as malignancy progresses the membrane capacitance decreases from approximately 26 mF/m2 to 15 mF/m2. This decline highlights the difference in dielectric properties and membrane composition between malignant and non-tumorous cells, reinforcing the potential of electrical characterization methods in distinguishing cancer progression.
@@
While dielectric properties provide a promising biophysical basis for distinguishing between malignant and non-malignant cells, the complexity and variability of the measured parameters necessitate robust computational tools for accurate classification. Artificial Intelligence (AI) and Machine Learning (ML) algorithms have been increasingly adopted to analyze bioelectrical data. These models can learn patterns from features such as membrane capacitance, conductivity, and impedance spectra to reliably differentiate between healthy and cancerous cells. Among various ML techniques, models such as Random Forest (RF), Support Vector Machine (SVM) and K-Nearest Neighbor (KNN), and Neural Network have shown promising results in biomedical classification tasks due to their ability to handle non-linear data and high dimension feature space [Breiman2001],[Liu2012],[Peterson2009],[Kramer2013],[Hearst1998],[Noble2006],[Hossain2021].